a loop of wire area = 0.09m^2 is perpendicular to a field of 3 tesla. if the loop rotates 1/4 turn in 0.1 sec, so it is parallel to the field, what emf is induced?
Answer #1

Area under magnetic field after (1/4) rev

A2 = 0 m^2

Emf induced = flux change / time

e = B ( A1 - 0) / t

e = 3* 0.09 / 0.1 = 2.7 V

Comment in case any doubt please rate my answer ...